OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (6, −3) and has a slope of 1/2. y − 6 = 1/2 (x + 3) y = x − 6 y + 3 = 1/2 (x − 6) x − 2y = 12

OpenStudy (aripotta):

y – y1 = m(x – x1) is point-slope form if y1 = -3, x1 = 6, and m = 1/2, what would our equation look like?

OpenStudy (anonymous):

i have no clue.

OpenStudy (aripotta):

all you have to do is plug in the numbers

OpenStudy (aripotta):